# Topology of Discrete Quantum Feedback Control

### v2

@@ Abstract
A general framework for analyzing topology of quantum channels of single-particle systems is developed to find a class of genuinely dynamical topological phases that can be realized by means of discrete quantum feedback control. We provide a symmetry classification of quantum channels by identifying ten symmetry classes of discrete quantum feedback control with projective measurements. We construct various types of topological feedback control by using topological Maxwell’s demons that achieve robust feedback-controlled chiral or helical transport against noise and decoherence. Topological feedback control thus offers a versatile tool for creating and controlling nonequilibrium topological phases in open quantum systems that are distinct from non-Hermitian and Lindbladian systems and should provide a guiding principle for topology-based design of quantum feedback control.
@@ I Introduction
Topology plays a key role in classification and characterization of diverse physical phenomena that are robust against local disturbances. In particular, topological phases of matter have been discussed in a wide variety of quantum and classical systems, which are characterized by global geometric structures of eigenstates rather than local order parameters [1, 2, 3, 4, 5, 6]. A quintessential example of topological quantum phenomena is the quantum Hall effect, where chiral transport immune to disorder emerges along the edges of a system [7, 8, 9, 10, 11]. Furthermore, topological phases are enriched by symmetry. For example, time-reversal symmetry permits helical edge states, where time-reversal partners flow in opposite directions with no net mass current, a phenomenon known as the quantum spin Hall effect [12, 13, 14, 15, 16]. Classification of phases of matter through integration of topology and symmetry has been of central interest in condensed matter physics [17, 18, 19].
@@
Feedback control provides a versatile tool for a wide range of applications in science and technology [40]. In particular, in the research arena of high-precision measurement and control of quantum systems, feedback control has achieved suppression of quantum noise [41], stabilization of desired states [42, 43], cooling [44, 45, 46], and quantum error correction [47, 48, 49, 50]. Feedback control also allows operations that would otherwise be impossible. A prime example is Maxwell’s demon, which was originally conceived as an intelligent being that can decrease entropy of a system against the second law of thermodynamics [51]. In modern parlance, Maxwell’s demon can be formulated as a feedback controller [52, 53], and various types of Maxwell’s demon have experimentally been realized in both classical [54, 55, 56, 57] and quantum [58, 59, 60, 61, 62, 63] systems. In view of versatile applicability of feedback control, we are naturally led to the following question: can we utilize feedback control to create and manipulate nonequilibrium topological phases? This is the problem we wish to address in this paper.
@@
Figure 1: (a) Schematic illustration of the eigenspectrum $\{\xi\}$ (shown by a blue curve and two blue dots) of a quantum channel for topological quantum feedback control. The spectrum involves a closed loop characterized by a winding number and the highly degenerate zero eigenvalues due to local measurement through which nonlocal coherence is projected out. The unit eigenvalue corresponds to a steady state. (b) Discrete quantum feedback control is comprised of a quantum measurement and a subsequent feedback operation that depends on the measurement outcome. A topologically nontrivial winding of the eigenspectrum of a quantum channel [see (a)] implies chiral (unidirectional) transport of a particle on a lattice due to feedback control.
Table 1: Comparison of different classes topological phases. The first row shows five classes of topological phases characterized by the topology of relevant operators shown in the second row. The third row indicates whether the dynamics is static, discrete, or continuous. The fourth row shows whether the system is closed or open. The fifth row shows whether the state of the system is pure or mixed.
@@
Closed/open Closed Closed Open Open Open
Pure/mixed Pure Pure Pure Mixed Mixed
@@
We develop a general framework for analyzing topology of quantum channels for discrete quantum feedback control. Mathematically, quantum channels are described by completely positive and trace-preserving (CPTP) maps, which provide a general form of time evolution in open quantum systems [47]. Here, the time evolution of a system under feedback control is nonunitary due to measurement backaction. Consequently, topology of CPTP maps is intrinsically non-Hermitian and beyond the theoretical framework of static (Hermitian) topological phases. We introduce topological invariants of quantum channels for single-particle systems on lattices under the general assumption of locality in measurement and feedback. For example, a nonzero winding number of the eigenspectrum of a quantum channel implies chiral transport due to feedback control, which may be regarded as a feedback-controlled counterpart of chiral edge transport in quantum Hall systems as schematically illustrated in Fig. 1.
@@
Quantum feedback control creates a class of nonequilibrium open quantum systems distinct from those described by non-Hermitian Hamiltonians [33, 34, 35] and Lindblad master equations [64, 65, 66, 67]. We summarize the comparison of different classes of topological phases in Table 1. In particular, discrete quantum feedback control concerns discrete time evolution from an initial quantum state to a post-feedback quantum state. Such time evolution cannot, in general, be expressed by using a generator of continuous time evolution. In fact, as shown in Sec. III, CPTP maps with local Kraus operators have highly degenerate zero eigenvalue, and therefore cannot be described by using any generator [see Eq. (39)]. Physically, the existence of zero eigenvalues can be understood as a consequence of local measurement through which nonlocal coherence in an extended system is projected out [see Fig. 1(a) and Sec. III]. A non-Hermitian Hamiltonian can only describe the dynamics of pure states and thus requires postselection of particular measurement outcomes [68]. The Lindblad quantum master equation can describe the dynamics of mixed states but does not refer to measurement outcomes [40]. In contrast, nonequilibrium topological phases created by quantum feedback control are realized by a nontrivial interplay between quantum measurement and feedback operations. We note that non-Hermitian topology has been applied to quantum channels in Ref. [33]. However, its application beyond zero-dimensional systems is nontrivial and has remained elusive; in fact, the locality of Kraus operators which is relevant in higher dimensions, is the key for defining topological invariants, as discussed in Sec. III.
@@
We present prototypical models of topological feedback control by using quantum Maxwell’s demons, which we shall refer to as topological Maxwell’s demons. In one dimension, a topological Maxwell’s demon causes feedback-assisted chiral transport due to topology. We find that the topological Maxwell’s demon creates topologically distinct phases by changing the duration of feedback control. We also show that feedback control can induce the non-Hermitian skin effect, which is a hallmark of dynamical topology unique to open systems [69, 70, 33, 71, 72, 73, 38, 74]. The chiral transport caused by the topological Maxwell’s demon is protected by topology and hence robust against disorder, noise, and decoherence during the feedback process.
@@
Symmetry plays a pivotal role in classification of topological phases [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we perform a symmetry classification of quantum feedback control. We find that only a subset of the 38 symmetry classes of non-Hermitian operators [75, 34, 35] is compatible with quantum channels with projective measurements. We identify ten relevant symmetry classes of quantum channels for discrete quantum feedback control (see Table 2) and provide examples of feedback control for each class. In particular, we propose a quantum Maxwell’s demon with a two-point measurement as an example of symmetry-protected topological feedback control. This demon generates helical spin transport protected by symmetry, which may be regarded as a counterpart of the quantum spin Hall effect in feedback-controlled quantum systems.
@@
Table 2: Ten-fold symmetry classification of discrete quantum feedback control with projective measurement. Each symmetry class is defined according to the presence or absence of $C,K$, and $Q$ symmetries in the CPTP map. Here, $C,K$, and $Q$ symmetries of a CPTP map $\mathcal{E}$ are defined by $\mathcal{V}_{C}\mathcal{E}^{T}\mathcal{V}_{C}^{-1}=\mathcal{E},\mathcal{V}_{K}\mathcal{E}^{*}\mathcal{V}_{K}^{-1}=\mathcal{E}$, and $\mathcal{V}_{Q}\mathcal{E}^{\dagger}\mathcal{V}_{Q}^{-1}=\mathcal{E}$, respectively, where $\mathcal{V}_{X}$ ($X=C,K,Q$) is a unitary superoperator. In the table, 00 indicates the absence of symmetry, and ±$\pm 1$ indicates the presence of symmetry with $\mathcal{V}_{X}\mathcal{V}_{X}^{*}=\pm\mathcal{I}$, where $\mathcal{I}$ denotes the identity superoperator. The labeling of the symmetry classes follows Ref. [34], where psH denotes pseudo-Hermiticity and the subscript $+$ (--) of psH indicates that the unitary superoperator for $C$ or $K$ symmetry commutes (anticommutes) with that of the pseudo-Hermiticity. In addition to the symmetry classes of a CPTP map $\mathcal{E}$, the corresponding Altland-Zirnbauer classes of $i\mathcal{E}$ are shown in the parentheses. In the fifth column, examples of conditions for feedback (FB) control in each symmetry class are shown (see Sec. VI for detailed discussions). CPTP maps with $C$ or $Q$ symmetry can be realized with feedback control with a two-point measurement (2PM). Symmetry classes without $K$ symmetry with $\mathcal{V}_{K}\mathcal{V}_{K}^{*}=+\mathcal{I}$ require unitary symmetry (US) due to the intrinsic modular conjugation symmetry of CPTP maps. AUS- denotes antiunitary symmetry whose square gives $-1$. Some symmetry classes may be realized with an additional feedback control that enforces the symmetry condition (see examples in Sec. VII). The rightmost column shows concrete models of topological feedback control, which are presented in Sec. IV and Sec. VII.
@@
AI (D) 00 $+1$ 00 No symmetry or particle-hole symmetry Chiral Maxwell’s demon / SSH demon
AI $+$ psH- (DIII) $-1$ $+1$ $1$ 2PM, reciprocity, and additional FB Helical Maxwell’s demon
@@
The rest of this paper is organized as follows. In Sec. II, we introduce quantum channels that describe discrete quantum feedback control. In Sec. III, we use the Bloch theory of translationally invariant quantum channels to show that topological invariants of quantum channels can be defined when Kraus operators are local in space. In Sec. IV, we construct an example of topological feedback control, which shows feedback-induced chiral transport due to its point-gap topology. In Sec. V, we examine symmetry of discrete quantum feedback control and identify ten symmetry classes that are compatible with projective measurement. We also discuss how to construct feedback control in each symmetry class in Sec. VI. On the basis of the symmetry classification, we provide various models of symmetry-protected topological feedback control in Sec. VII. In particular, we present topological feedback control protected by transposition symmetry which exhibits helical spin transport due to an interplay between symmetry and topology. In Sec. VIII, we discuss topological phenomena arising from complete positivity and a trace-preserving property, which are unique to quantum channels and absent in general non-Hermitian systems. In Sec. IX, we extend our theory to the case of nonunitary feedback control, thereby showing robustness of topological feedback control against dissipation. We also propose feedback control with engineered dissipation, where dissipation is used as an essential tool to achieve feedback control. We conclude this paper in Sec. X with an outlook. Some technical details are relegated to the Appendices to avoid digressing from the main subject. In Appendix A, we present the detail of the Bloch theory of quantum channels. In Appendix B, we describe the Bloch theory of quantum feedback control with projective measurement. In Appendix C, we discuss a spectral property of quantum channels for projective measurement. In Appendix D, we summarize the correspondence between symmetry of time-evolution operators and that of their generators. In Appendix E, we provide a supplementary discussion of symmetry-preserving boundary conditions for a $\mathbb{Z}_{2}$ Maxwell’s demon. In Appendix F, we present a model of topological feedback control that shows a non-Hermitian skin effect but with no boundary-localized steady state.
@@ II Discrete quantum feedback control
where the index $m$ specifies a measurement outcome and $\hat{I}$ is the identity operator [40]. The probability of an outcome $m$ is given by $p_{m}=\mathrm{Tr}[\hat{M}_{m}^{\dagger}\hat{M}_{m}\hat{\rho}_{\mathrm{i}}]$, and the postmeasurement state corresponding to an outcome $m$ is given by
@@
where the nonunitary state change is caused by measurement backaction. Then, we let the system undergo unitary time evolution during time $\tau$, which is described by a unitary operator
@@
where $\hat{H}_{m}(t)$ is the feedback Hamiltonian which depends on the measurement outcome $m$ and $\mathcal{T}$ is the time-ordering operator. We set $\hbar=1$ throughout this paper. If we take an ensemble average over measurement outcomes, then the final density matrix after one feedback loop is given by [52]
@@
which is a completely positive and trace-preserving (CPTP) map [47], where $\hat{K}_{m}:=\hat{U}_{m}\hat{M}_{m}$ is the Kraus operator satisfying $\sum_{m}\hat{K}_{m}^{\dagger}\hat{K}_{m}=\hat{I}$.
If a feedback cycle is repeated, the final density matrix is given by applying CPTP maps multiple times. Suppose that we repeat a measurement and a feedback operation $N$ times. Let $(m_{1},m_{2},\cdots,m_{N})$ be a sequence of measurement outcomes through this process. The density matrix after the $n$th measurement is given by
@@
The above description can be generalized to the cases in which different measurement and feedback operations are performed multiple times during a single cycle. For example, suppose that we perform a second measurement described by measurement operators $\{\hat{M}_{m_{2}}^{\prime}\}$ after the first measurement and the subsequent feedback operation. Then, we perform the second feedback control with unitary operators $\{\hat{U}_{m_{1},m_{2}}^{\prime}\}$, which may depend on the outcomes of the first and second measurements. The overall feedback cycle is described by a CPTP map
@@
where $\xi_{n}\in\mathbb{C}$ denotes an eigenvalue, and $\hat{\rho}_{n}^{R}$ and $\hat{\rho}_{n}^{L}$ are the corresponding right and left eigenoperators which satisfy the biorthogonal relation $\mathrm{Tr}[(\hat{\rho}_{m}^{L})^{\dagger}\hat{\rho}_{n}^{R}]=0$ for $m\neq n$. In terms of the eigensystem of the CPTP map $\mathcal{E}$, the density matrix after a feedback loop in Eq. (4) is expanded as
@@
is an expansion coefficient which depends on the initial state $\hat{\rho}_{\mathrm{i}}$. Here, for simplicity we assume that the CPTP map is diagonalizable. If the CPTP map is not diagonalizable, we can expand the density matrix by using generalized eigenmodes [68]. Thus, the eigenvalues and the (generalized) eigenmodes of the CPTP map completely characterize the discrete quantum feedback control.
The eigensystem of a CPTP map has some general properties (see Fig. 1(a) for a schematic illustration). First, all the eigenvalues $\xi_{n}$ of a CPTP map satisfies $|\xi_{n}|\leq 1$ because of the contraction property of a CPTP map [76]. An eigenmode $\hat{\rho}_{\mathrm{SS}}$ with eigenvalue $+1$ is a steady state of the dynamics in the sense that $\mathcal{E}(\hat{\rho}_{\mathrm{SS}})=\hat{\rho}_{\mathrm{SS}}$. Second, since a CPTP map preserves the trace of the density matrix, that is, $\mathrm{Tr}[\mathcal{E}(\hat{\rho})]=\mathrm{Tr}[\hat{\rho}]$, all the right eigenmodes except for those corresponding to steady states are traceless: $\mathrm{Tr}[\hat{\rho}_{n}^{R}]=0\ (\xi_{n}\neq 1)$. Third, a CPTP map preserves the Hermiticity of an operator since
@@
It follows from Eqs. (12) and (16) that if $\hat{\rho}_{n}^{R}$ is an eigenmode of $\mathcal{E}$ with eigenvalue $\xi_{n}$, then $(\hat{\rho}_{n}^{R})^{\dagger}$ is also an eigenmode with eigenvalue $\xi_{n}^{*}$ since

### v3

@@ Abstract
A general framework for analyzing the topology of quantum channels of single-particle systems is developed to find a class of genuinely dynamical topological phases that can be realized by means of discrete quantum feedback control. We provide a symmetry classification of quantum channels by identifying ten symmetry classes of discrete quantum feedback control with projective measurements. We construct various types of topological feedback control by using topological Maxwell demons that achieve robust feedback-controlled chiral or helical transport against noise and decoherence. Topological feedback control thus offers a versatile tool for creating and controlling nonequilibrium topological phases in open quantum systems that are distinct from non-Hermitian and Lindbladian systems and should provide a guiding principle for topology-based design of quantum feedback control.
@@ I Introduction
Topology plays a key role in the classification and characterization of diverse physical phenomena that are robust against local disturbances. In particular, topological phases of matter have been discussed in a wide variety of quantum and classical systems, which are characterized by global geometric structures of eigenstates rather than local order parameters [1, 2, 3, 4, 5, 6]. A quintessential example of topological quantum phenomena is the quantum Hall effect, where chiral transport immune to disorder emerges along the edges of a system [7, 8, 9, 10, 11]. Furthermore, topological phases are enriched by symmetry. For example, time-reversal symmetry permits helical edge states, where time-reversal partners flow in opposite directions with no net mass current, a phenomenon known as the quantum spin Hall effect [12, 13, 14, 15, 16]. Classification of phases of matter through integration of topology and symmetry has been of central interest in condensed matter physics [17, 18, 19].
@@
Feedback control provides a versatile tool for a wide range of applications in science and technology [40]. In particular, in the research arena of high-precision measurement and control of quantum systems, feedback control has achieved suppression of quantum noise [41], stabilization of desired states [42, 43], cooling [44, 45, 46], and quantum error correction [47, 48, 49, 50]. Feedback control also allows operations that would otherwise be impossible. A prime example is Maxwell’s demon, which was originally conceived as an intelligent being that can decrease entropy of a system against the second law of thermodynamics [51]. Thus, Maxwell’s demon can be formulated as a feedback controller [52, 53], and various types of Maxwell demons have experimentally been realized in both classical [54, 55, 56, 57] and quantum [58, 59, 60, 61, 62, 63] systems. In view of versatile applicability of feedback control, we are naturally led to the following question: Can we utilize feedback control to create and manipulate nonequilibrium topological phases? We address this problem in the following.
@@
Figure 1: (a) Schematic illustration of the eigenspectrum $\{\xi\}$ (shown by a blue curve and two blue dots) of a quantum channel for topological quantum feedback control. The spectrum involves a closed loop characterized by a winding number and the highly degenerate zero eigenvalues due to local measurement through which nonlocal coherence is projected out. The unit eigenvalue corresponds to a steady state. (b) Discrete quantum feedback control comprised of a quantum measurement and a subsequent feedback operation that depends on the measurement outcome. A topologically nontrivial winding of the eigenspectrum of a quantum channel [see diagram (a)] implies chiral (unidirectional) transport of a particle on a lattice due to feedback control.
Table 1: Comparison of different classes of topological phases. The first row shows five classes of topological phases characterized by the topology of relevant operators shown in the second row. The third row indicates whether the dynamics is static, discrete, or continuous. The fourth row shows whether the system is closed or open. The fifth row shows whether the state of the system is pure or mixed.
@@
Closed or open Closed Closed Open Open Open
Pure or mixed Pure Pure Pure Mixed Mixed
@@
We develop a general framework for analyzing the topology of quantum channels for discrete quantum feedback control. Mathematically, quantum channels are described by completely positive and trace-preserving (CPTP) maps, which provide a general form of time evolution in open quantum systems [47]. Here, the time evolution of a system under feedback control is nonunitary due to measurement backaction. Consequently, the topology of CPTP maps is intrinsically non-Hermitian and beyond the theoretical framework of static (Hermitian) topological phases. We introduce topological invariants of quantum channels for single-particle systems on lattices under the general assumption of locality in measurement and feedback. For example, a nonzero winding number of the eigenspectrum of a quantum channel implies chiral transport due to feedback control, which may be regarded as a feedback-controlled counterpart of chiral edge transport in quantum Hall systems as schematically illustrated in Fig. 1.
@@
Quantum feedback control creates a class of nonequilibrium open quantum systems distinct from those described by non-Hermitian Hamiltonians [33, 34, 35] and Lindblad master equations [64, 65, 66, 67]. We summarize the comparison of different classes of topological phases in Table 1. In particular, discrete quantum feedback control concerns discrete time evolution from an initial quantum state to a postfeedback quantum state. Such time evolution cannot, in general, be expressed by using a generator of continuous time evolution. In fact, as shown in Sec. III, CPTP maps with local Kraus operators have highly degenerate zero eigenvalues, and therefore cannot be described by using any generator [see Eq. (39)]. Physically, the existence of zero eigenvalues can be understood as a consequence of local measurement through which nonlocal coherence in the system is projected out [see Fig. 1(a) and Sec. III]. A non-Hermitian Hamiltonian can only describe the dynamics of pure states and requires postselection of particular measurement outcomes [68]. The Lindblad quantum master equation can describe the dynamics of mixed states but does not refer to measurement outcomes [40]. In contrast, nonequilibrium topological phases created by quantum feedback control are realized by a nontrivial interplay between quantum measurement and feedback operations. We note that non-Hermitian topology has been applied to quantum channels in Ref. [33]. However, its application beyond zero-dimensional systems is nontrivial and has remained elusive; in fact, the locality of Kraus operators, which is relevant in higher dimensions, is key for defining topological invariants, as discussed in Sec. III.
@@
We present prototypical models of topological feedback control by using quantum Maxwell demons, which we shall refer to as topological Maxwell demons. In one dimension, a topological Maxwell demon causes feedback-assisted chiral transport due to topology. We find that the topological Maxwell demon creates topologically distinct phases by changing the duration of feedback control. We also show that feedback control can induce the non-Hermitian skin effect, which is a hallmark of dynamical topology unique to open systems [69, 70, 33, 71, 72, 73, 38, 74]. The chiral transport caused by the topological Maxwell demon is protected by topology and hence robust against disorder, noise, and decoherence during the feedback process.
@@
Symmetry plays a pivotal role in the classification of topological phases [18, 17, 19, 27, 28, 33, 34, 35, 65, 37, 66, 67]. Here, we perform a symmetry classification of quantum feedback control. We find that only a subset of the 38 symmetry classes of non-Hermitian operators [75, 34, 35] is compatible with quantum channels with projective measurements. We identify ten relevant symmetry classes of quantum channels for discrete quantum feedback control (see Table 2) and provide examples of feedback control for each class. In particular, we propose a quantum Maxwell demon with a two-point measurement as an example of symmetry-protected topological feedback control. This demon generates helical spin transport protected by symmetry, which may be regarded as a counterpart of the quantum spin Hall effect in feedback-controlled quantum systems.
@@
Table 2: Tenfold symmetry classification of discrete quantum feedback control with projective measurements. Each symmetry class is defined according to the presence or absence of $C,K$, and $Q$ symmetries in the CPTP map. Here, $C,K$, and $Q$ symmetries of a CPTP map $\mathcal{E}$ are defined by $\mathcal{V}_{C}\mathcal{E}^{T}\mathcal{V}_{C}^{-1}=\mathcal{E},\mathcal{V}_{K}\mathcal{E}^{*}\mathcal{V}_{K}^{-1}=\mathcal{E}$, and $\mathcal{V}_{Q}\mathcal{E}^{\dagger}\mathcal{V}_{Q}^{-1}=\mathcal{E}$, respectively, where $\mathcal{V}_{X}$ ($X=C,K,Q$) is a unitary superoperator. Here, 00 indicates the absence of symmetry, and ±$\pm 1$ indicates the presence of symmetry with $\mathcal{V}_{X}\mathcal{V}_{X}^{*}=\pm\mathcal{I}$, where $\mathcal{I}$ denotes the identity superoperator. The labeling of the symmetry classes follows Ref. [34], where psH denotes pseudo-Hermiticity and the subscript $+$ (--) of psH indicates that the unitary superoperator for $C$ or $K$ symmetry commutes (anticommutes) with that of the pseudo-Hermiticity. In addition to the symmetry classes of a CPTP map $\mathcal{E}$, the corresponding Altland-Zirnbauer classes of $i\mathcal{E}$ are shown in the parentheses. In the fifth column, examples of conditions for feedback (FB) control in each symmetry class are shown (see Sec. VI for detailed discussions). CPTP maps with $C$ or $Q$ symmetry can be realized with feedback control with a two-point measurement (2PM). Symmetry classes without $K$ symmetry with $\mathcal{V}_{K}\mathcal{V}_{K}^{*}=+\mathcal{I}$ require unitary symmetry (US) due to the intrinsic modular conjugation symmetry of CPTP maps. AUS- denotes antiunitary symmetry whose square gives $-1$. Some symmetry classes may be realized with an additional feedback control that enforces the symmetry condition (see examples in Sec. VII). The rightmost column shows concrete models of topological feedback control, which are presented in Secs. IV and VII.
@@
AI (D) 00 $+1$ 00 No symmetry or particle-hole symmetry Chiral Maxwell demon / SSH demon
AI $+$ psH- (DIII) $-1$ $+1$ $1$ 2PM, reciprocity, and additional FB Helical Maxwell demon
@@
The rest of this paper is organized as follows. In Sec. II, we introduce quantum channels that describe discrete quantum feedback control. In Sec. III, we use the Bloch theory of translationally invariant quantum channels to show that topological invariants of quantum channels can be defined when Kraus operators are local in space. In Sec. IV, we construct an example of topological feedback control, which shows feedback-induced chiral transport due to its point-gap topology. In Sec. V, we examine the symmetry of discrete quantum feedback control and identify ten symmetry classes that are compatible with projective measurements. We also discuss how to construct feedback control in each symmetry class in Sec. VI. On the basis of the symmetry classification, we provide various models for symmetry-protected topological feedback control in Sec. VII. In particular, we present topological feedback control protected by transposition symmetry, which exhibits helical spin transport due to an interplay between symmetry and topology. In Sec. VIII, we discuss topological phenomena arising from complete positivity and a trace-preserving property, which are unique to quantum channels and absent in general non-Hermitian systems. In Sec. IX, we extend our theory to the case of nonunitary feedback control, thereby showing the robustness of topological feedback control against dissipation. We also propose feedback control with engineered dissipation, where dissipation is used as an essential ingredient to achieve feedback control. We conclude this paper in Sec. X with an outlook. Some technical details are relegated to the appendixes to avoid digressing from the main subject. In Appendix A, we present the details of the Bloch theory of quantum channels. In Appendix B, we describe the Bloch theory of quantum feedback control with projective measurements. In Appendix C, we discuss a spectral property of quantum channels for projective measurements. In Appendix D, we summarize the correspondence between symmetry of time-evolution operators and that of their generators. In Appendix E, we provide a supplementary discussion of symmetry-preserving boundary conditions for a $\mathbb{Z}_{2}$ Maxwell demon. In Appendix F, we present a model of topological feedback control that shows a non-Hermitian skin effect but no boundary-localized steady state.
@@ II Discrete quantum feedback control
where the index $m$ specifies a measurement outcome and $\hat{I}$ is the identity operator [40]. The probability of an outcome $m$ being obtained is given by $p_{m}=\mathrm{Tr}[\hat{M}_{m}^{\dagger}\hat{M}_{m}\hat{\rho}_{\mathrm{i}}]$, and the postmeasurement state corresponding to an outcome $m$ is given by
@@
where the nonunitary state change is caused by measurement backaction. Next, we let the system undergo unitary time evolution during time $\tau$, which is described by a unitary operator
@@
where $\hat{H}_{m}(t)$ is the feedback Hamiltonian that depends on the measurement outcome $m$ and $\mathcal{T}$ is the time-ordering operator. We set $\hbar=1$ throughout this paper. If we take an ensemble average over measurement outcomes, then the final density matrix after one feedback loop is given by [52]
@@
which is a CPTP map [47], where $\hat{K}_{m}:=\hat{U}_{m}\hat{M}_{m}$ is the Kraus operator satisfying $\sum_{m}\hat{K}_{m}^{\dagger}\hat{K}_{m}=\hat{I}$.
If a feedback cycle is repeated, the final density matrix is given by applying CPTP maps multiple times. Suppose that we repeat a measurement and a feedback operation $N$ times. Let $(m_{1},m_{2},\cdots,m_{N})$ be a sequence of measurement outcomes in this process. The density matrix after the $n$th measurement is given by
@@
The above description can be generalized to the cases in which different measurements and feedback operations are performed during a single cycle. For example, suppose that we perform a second measurement described by measurement operators $\{\hat{M}_{m_{2}}^{\prime}\}$ after the first measurement and the subsequent feedback operation. Then, we perform the second feedback control with unitary operators $\{\hat{U}_{m_{1},m_{2}}^{\prime}\}$, which may depend on the outcomes of the first and second measurements. The entire feedback cycle is described by a CPTP map
@@
where $\xi_{n}\in\mathbb{C}$ denotes an eigenvalue, and $\hat{\rho}_{n}^{R}$ and $\hat{\rho}_{n}^{L}$ are the corresponding right and left eigenoperators that satisfy the biorthogonal relation $\mathrm{Tr}[(\hat{\rho}_{m}^{L})^{\dagger}\hat{\rho}_{n}^{R}]=0$ for $m\neq n$. In terms of the eigensystem of the CPTP map $\mathcal{E}$, the density matrix after a feedback loop in Eq. (4) is expanded as
@@
is an expansion coefficient that depends on the initial state $\hat{\rho}_{\mathrm{i}}$. Here, for simplicity, we assume that the CPTP map is diagonalizable. If the CPTP map is not diagonalizable, we can expand the density matrix by using generalized eigenmodes [68]. Thus, the eigenvalues and the (generalized) eigenmodes of the CPTP map completely characterize the discrete quantum feedback control.
The eigensystem of a CPTP map has some general properties [see Fig. 1(a) for a schematic illustration]. First, all the eigenvalues $\xi_{n}$ of a CPTP map satisfy $|\xi_{n}|\leq 1$ because of the contraction property of a CPTP map [76]. An eigenmode $\hat{\rho}_{\mathrm{SS}}$ with eigenvalue $+1$ is a steady state of the dynamics in the sense that $\mathcal{E}(\hat{\rho}_{\mathrm{SS}})=\hat{\rho}_{\mathrm{SS}}$. Second, since a CPTP map preserves the trace of the density matrix, that is, $\mathrm{Tr}[\mathcal{E}(\hat{\rho})]=\mathrm{Tr}[\hat{\rho}]$, all the right eigenmodes except for those corresponding to steady states are traceless: $\mathrm{Tr}[\hat{\rho}_{n}^{R}]=0\ (\xi_{n}\neq 1)$. Third, a CPTP map preserves the Hermiticity of an operator since
@@
It follows from Eqs. (12) and (16) that if $\hat{\rho}_{n}^{R}$ is an eigenmode of $\mathcal{E}$ with eigenvalue $\xi_{n}$, then $(\hat{\rho}_{n}^{R})^{\dagger}$ is an eigenmode with eigenvalue $\xi_{n}^{*}$ since